# Observational constraints on degeneracy of non-canonical inflation driven by quartic potential

### v2

@@ III Reheating epoch
As mentioned before, at the end of inflation, the kinetic energy of the inflaton overcomes its potential energy. Thence the inflaton oscillates around the minimum of the potential and gradually releases its energy. This epoch is called reheating and its nature is still not fully understood Baumann:2009 . Considering the reheating epoch can also be useful for verifying inflationary models. In this epoch, the equation of state parameter ($\omega_{re}$), the duration of the reheating epoch ($N_{re}$), and the reheating temperature ($T_{re}$) are usually investigated. The duration of reheating epoch and reheating temperature both depend on the equation of state parameter. The most important constraint that the duration of the reheating epoch imposes on the model is that its value must be positive, i.e., $N_{re}>0$. Additionally, the duration of the reheating epoch should be significantly shorter than the duration of the inflation epoch and it can be obtained as follows Martin2015 ; Munoz2015 ; Cook2015
$N_{re}=\left(\frac{4}{1-3\omega_{re}}\right)\left[-\frac{1}{4}\ln\left(\frac{30}{\pi^{2}g_{re}}\right)-\frac{1}{3}\ln\left(\frac{11g_{re}}{43}\right)-\ln\left(\frac{k_{\ast}}{a_{0}T_{0}}\right)-\ln\left(\frac{\rho_{e}^{1/4}}{H_{\ast}}\right)-N\right],$ (41)
where $g_{re}=106.75$ is the effective number of relativistic degrees of freedom for energy during the reheating epoch, $k_{\ast}=0.05~{}{\rm Mpc^{-1}}$ is the pivot scale, $a_{0}=1$ is the scale factor at the present time, $T_{0}=2.725~{}{\rm K}$ is the current temperature of CMB radiation, $\rho_{e}$ is the energy density of the inflaton field at the end of inflation, $H_{\ast}$ is the Hubble parameter at the horizon crossing time and $N$ defines the duration of inflation from the end of inflation until the horizon passing moment . Using the conversion factor $1~{}{\rm Mpc^{-1}}=6.39\times 10^{-39}{\rm GeV}$, Eq. (41) can be simplified as follows
$N_{re}=\frac{4}{1-3\omega_{re}}\left[61.6-\ln\left(\frac{\rho_{e}^{1/4}}{H_{\ast}}\right)-N\right],$ (42)
@@
It is known that, the equation of state parameter is defined as
$\omega_{\phi}\equiv\frac{P_{\phi}}{\rho_{\phi}}=-1-\frac{2}{3}\frac{\dot{H}}{H^{2}}=-1+\frac{2}{3}\epsilon,$ (44)
where $\epsilon=1$ at the end of inflation. Therefore, the value of the equation of state parameter at the end of inflation is $\omega_{\phi_{e}}=-\frac{1}{3}$, and consequently
$P_{e}=-\frac{1}{3}\rho_{e}.$ (45)
By substituting Eqs. (3) and (4) into Eq. (45), the energy density at the end of inflation $\rho_{e}$ in the non-canonical framework with power-law Lagrangian (2) can be obtained by
$\rho_{e}=\left(\frac{3\alpha}{\alpha-1}\right)V_{e},$ (46)
@@
Note that Eq. (8) has a singularity at $\dot{\phi}\rightarrow 0$. According to the calculations of Ref. Unnikrishnan:2012 , this singularity in the equation of motion of the non-canonical model with power-law Lagrangian (2), can be removed by the following regularized Lagrangian
${\cal L}_{R}=\left(\frac{X}{1+\beta}\right)\left(1+\beta\left(\frac{X}{M^{4}}\right)^{\alpha-1}\right)-V\left(\phi\right),$ (47)
where $\beta$ is a dimensionless parameter. In the limit $\beta\gg 1$, Eq. (8) becomes
$\ddot{\phi}+\frac{3H\dot{\phi}}{2\alpha-1}+\left(\frac{{V}^{\prime}\left(\phi\right)}{\epsilon+\alpha\left(2\alpha-1\right)\left(\frac{X}{M^{4}}\right)^{\alpha-1}}\right)=0,~{}~{}X=\frac{1}{2}\dot{\phi}^{2},$ (48)
wherein $\epsilon\equiv\left(1+\beta\right)^{-1}$. Since $\omega_{\phi}\equiv\frac{p_{\phi}}{\rho_{\phi}}$, for $\beta\gg 1$ we have
$1+\left\langle\omega_{\phi}\right\rangle=\left(\frac{2\alpha}{M^{4\left(\alpha-1\right)}}\right)\left\langle\frac{X^{\alpha}}{\rho_{\phi}}\right\rangle,$ (49)
[…]
@@
Similar to the standard inflation in the canonical framework, we assume that the timescale of the scalar field oscillation is much smaller than that of the expansion of the universe. In this limit, the time variations of $\rho_{\phi}$ during each cycle is small enough to approximate $\rho_{\phi}\simeq V(\phi_{m})$, where $\phi_{m}$ is the maximum value of the field during a given cycle. According to Eq. (50), in the framework of non-canonical inflation with a power-law Lagrangian for the potential $V\left(\phi\right)=V_{0}\phi^{n}$, the equation of state parameter can be calculated as follows
$\omega_{re}=\left\langle\omega_{\phi}\right\rangle=\frac{n-2\alpha}{n\left(2\alpha-1\right)+2\alpha},$ (51)
@@
$\omega_{re}=\frac{2-\alpha}{5\alpha-2}.$ (52)
In the following, we calculate the duration of the reheating epoch $N_{re}$ for non-canonical action with power-law Lagrangian (2) and quartic potential (35) by inserting Eqs. (43), (46) and (52) in (42). Then, considering the relation between $n_{s}$ and $N$ from Eq. (36), we plot the values of $N_{re}$ in terms of the scalar spectral index $n_{s}$ in Fig. 4(a) for different values of $\alpha$ parameter according to Table 1. The red curves in this graph, corresponds to $\alpha=130$ which is obtained from the non-Gaussianity constraint. Also, the dark blue, yellow, dashed blue and dashed black curves represent $\alpha=8.6$, $\alpha=11.3$, $\alpha=18.8$, and $\alpha=110$, respectively. Furthermore, the dark (light) blue area indicates the 68% (95%) CL for the scalar spectral index, according to recent data of Planck 2018. It can be observed that despite variations in $\alpha$, there is little difference in the values obtained for the duration of reheating $N_{re}$. In other words, $N_{re}$ is not proportional to changes in $\alpha$. Therefore, in non-canonical framework with quartic potential, $N_{re}$ cannot be a suitable tool to break the degeneracy of the model with respect to $\alpha$ parameter.
Fig. 4(b) shows the length of the reheating epoch ($N_{re}$) versus the length of the inflation period ($N$) for different values of $\alpha$ parameter. As it can be seen from this figure, the length of the reheating epoch $N_{re}$ becomes negative at large $e$-folds. The condition $N_{re}\geq 0$ constrains the length of inflationary era in the way that it must be $N\leq 55.7$. On the other side, the diagram of Fig. 3 shows that the number of $e$-folds should be more than 55 as $N\geq 55$ to ensure the consistency of the model with the observational data of Planck 2018. Altogether, the duration of the inflationary era for this model is constrained as $55\leq N\leq 55.7$ from the observational data on $r-n_{s}$ diagram thereto the reheating consideration $N_{re}\geq 0$. Subsequently, the $\alpha$ parameter is more constrained too. In Table 2, the computed ranges of $\alpha$, $\mu$ and $N_{re}$ are shown for permitted durations of the inflationary era $55\leq N\leq 55.7$.
Table 2: The allowed ranges of the non-canonical parameters of the model ($\alpha$ and $\mu$) according to the constraints of $c_{s}^{2}+r+n_{s}+f_{\rm NL}^{\rm equil}+N_{re}$ for the permitted $e$-folds number $55\leq N\leq 55.7$.
  $N$   $\alpha$   $\mu~{}(\times 10^{-5})$   $N_{re}$
  $55$   $110\leq\alpha\leq 130$   $2.23\leq\mu\leq 2.30$   $1.10\leq N_{re}\leq 1.26$
  $55.3$   $80\leq\alpha\leq 130$   $2.22\leq\mu\leq 2.40$   $0.34\leq N_{re}\leq 0.81$
  $55.7$   $60\leq\alpha\leq 65$   $2.46\leq\mu\leq 2.49$   $0\leq N_{re}\leq 0.08$
Figure 4: Duration of the reheating stage $N_{re}$ $(a)$ versus the scalar spectral index $n_{s}$, $(b)$ versus the $e$-fold number $N$, for different values of $\alpha$ parameter according to Table 1. The dark and light blue regions show, respectively, the 68% and 95% CL of Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO data.
@@
In addition to $N_{re}$, the reheating temperature $T_{re}$ can also be used to survey inflationary models and it can be calculated as follows Martin2015 ; Munoz2015 ; Cook2015
$T_{re}=\left(\frac{45}{\pi^{2}g_{re}}\right)^{1/4}V_{e}^{1/4}\exp\left[-\frac{3}{4}(1+\omega_{re})N_{re}\right].$ (53)
@@
This parameter depends on $\omega_{re}$ and $N_{re}$. Since the reheating stage occurs after the end of inflation and before the beginning of the Big Bang Nucleosynthesis (BBN) stage, the upper and lower limits on the reheating temperature are defined as follows Martin2015
$1\,{\rm MeV}\leq T_{re}\leq 1.6\times 10^{16}\,{\rm GeV}.$ (54)
In Fig. 5, the reheating temperature $T_{re}$, Eq. (53), is plotted as a function of the scalar spectral index $n_{s}$ for different values of $\alpha$ parameter. The dark (light) blue area represents the 68% (95%) CL specified by Eq. (18) for the scalar spectral index according to recent data of Planck 2018. The gray area at the bottom of the diagram represents the constraint from BBN. The horizontal line at the top of the diagram indicates the upper limit of the reheating temperature. It is worth noting that changing in the $\alpha$ parameter does not yield the significant change in the reheating temperature. Therefore, the reheating temperature in this model, is degenerate with respect to changes in $\alpha$ parameter. It can be inferred from Figs. 4 and 5 that, the duration $N_{re}$ and the temperature $T_{re}$ of the reheating stage can impose constraint on $\alpha$ from free parameters of the inflationary model with a quartic potential in the non-canonical framework.
@@
Figure 5: The reheating temperature $T_{re}$ versus $n_{s}$ for different values of $\alpha$ parameter according to Table 2. The bottom gray hatched area corresponds to the BBN constraint. The maximum allowable reheating temperature $T_{re}=1.6\times 10^{16}~{}\rm GeV$ is indicated by a horizontal dashed black line at the top of the graph. The dark (light) blue regions shows, the 68% (95%) CL of Planck 2018 TT, TE, EE + LowE + Lensing + BK18 + BAO data.
@@ IV The relic gravitational waves
Inflation theory predicts that, the primordial tensor fluctuations of the inflaton field can lead to production of gravitational waves (GWs). During inflation, these fluctuations extend beyond the Hubble horizon, and upon reentering, they generate relic GWs Starobinsky:1979 ; allen88 ; sahni:1990 ; sami2002 ; dany_18 ; dany_19 ; Bernal:2019lpc . Due to their weak interaction with matter and radiation, these relic GWs serve as a valuable tool for probing the early universe. The present frequency of the originated relic GWs from the cosmic inflation can be calculated in terms of temperature using the following equation sahni:1990 ; dany_19 ; sami2002
@@
where $g_{0}^{s}=3.94$ and $g_{T}^{s}=106.75$ are the effective number of relativistic degrees of freedom in entropy at the present time and each epoch with temperature $T$, respectively. Furthermore, $g_{T}=106.75$ is the effective number of relativistic degrees of freedom in energy. The allowed ranges for the reheating temperature $T_{re}$ and the corresponding frequency of GWs in the reheating epoch $f_{re}$ are listed in Table 3 for several values of $\alpha$ parameter according to Table 2. The present density parameter spectrum of relic GWs in the radiation-dominated and reheating periods are calculated from the following two equations (for more information, see Ref. sahni:1990 ; dany_19 ; sami2002 )
@@
where $\Omega_{r_{0}}=2.47\times 10^{-5}h^{-2}$ is the radiation density parameter at present time, $f_{*}=2.4\times 10^{-16}~{}({\rm Hz})$ is the frequency of GWs at the CMB pivot scale $k_{\ast}=0.05~{}{\rm Mpc^{-1}}$, $f_{eq}=1.7\times 10^{-17}~{}({\rm Hz})$ represents the frequency of GWs at the matter-radiation equality and $f_{e}=4.3\times 10^{8}~{}({\rm Hz})$ is the frequency of GWs at the end of inflation.
Table 3: The allowed regions of the reheating temperature $T_{re}$ and the corresponding frequency $f_{re}$ of GWs in the reheating epoch computed by Eq. (55) for several values of $\alpha$ according to Fig. 5.
  $\alpha$   $T_{re}~{}(\times 10^{15}~{}{\rm GeV})$   $f_{re}~{}(\times 10^{7}~{}{\rm Hz})$
  $60$   $0.472\leq T_{re}\leq 1.432$   $1.261\leq f_{re}\leq 3.823$
  $80$   $0.542\leq T_{re}\leq 1.385$   $1.448\leq f_{re}\leq 3.697$
  $110$   $0.629\leq T_{re}\leq 1.334$   $1.680\leq f_{re}\leq 3.561$
  $130$   $0.680\leq T_{re}\leq 1.308$   $1.814\leq f_{re}\leq 3.492$
In Fig. 6, the present density parameter spectra of GWs are plotted, for the scalar spectral index $n_{s}=0.9583$, using Eqs. (55)-(57). The green area represents the sensitivity range of the Big Bang Observer (BBO) gravitational wave detector Yagi:2011 ; BBO:2003 , while the red area represents the sensitivity range of the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) detector Yagi:2011 ; Seto:2001 . The black vertical lines in the graph indicate the frequencies of GWs calculated for specific reheating temperatures according to Table 3. The blue and red lines in this graph represent the present energy density of GWs computed for the quartic potential in the non-canonical framework for $\alpha=80$ and $\alpha=130$, respectively. It can be observed that, the curves fall within the sensitivity range of GWs detectors. Hence, with the advancement of these detectors, there is a possibility of detecting the predicted relic GWs by this model. Moreover, varying in $\alpha$ parameter leads to distinct values of relic GWs density. Therefore, the analysis of the relic GWs can break the degeneracies of the predictions of the non-canonical inflationary model, containing the quartic potential (35) and power-law Lagrangian (2), with respect to the $\alpha$ parameter.
Figure 6: Present density parameter spectra of GWs versus frequency for $\alpha=80$ (blue curve) and $\alpha=130$ (red curve), considering the scalar spectral index $n_{s}=0.9583$. The hatched regions represent the sensitivity ranges of BBO and DECIGO GWs detectors. The vertical black lines indicate frequencies of GWs calculated for the indicated reheating temperatures.

### v3

@@ III Reheating epoch
As mentioned before, at the end of inflation, the kinetic energy of the inflaton overcomes its potential energy. Thence the inflaton oscillates around the minimum of the potential and gradually releases its energy. This epoch is called reheating and its nature is still not fully understood Baumann:2009 . Considering the reheating epoch can also be useful for verifying inflationary models. In this epoch, the equation of state parameter ($\omega_{\rm re}$), the duration of the reheating epoch ($N_{\rm re}$), and the reheating temperature ($T_{\rm re}$) are usually investigated. The duration of reheating epoch and reheating temperature both depend on the equation of state parameter. The most important constraint that the duration of the reheating epoch imposes on the model is that its value must be positive, i.e., $N_{\rm re}\geq 0$. Additionally, the duration of the reheating epoch should be significantly shorter than the duration of the inflation epoch and it can be obtained as follows Martin2015 ; Munoz2015 ; Cook2015 ; German2020
$N_{\rm re}=\left(\frac{4}{1-3\omega_{\rm re}}\right)\left[-\frac{1}{4}\ln\left(\frac{30}{\pi^{2}g_{\rm re}}\right)-\frac{1}{3}\ln\left(\frac{11g_{\rm re}}{43}\right)-\ln\left(\frac{k_{\ast}}{a_{0}T_{0}}\right)-\ln\left(\frac{\rho_{e}^{1/4}}{H_{\ast}}\right)-N\right],$ (41)
where $g_{\rm re}=106.75$ is the effective number of relativistic degrees of freedom for energy during the reheating epoch. Also $k_{\ast}=0.05~{}{\rm Mpc^{-1}}$ is the pivot scale, $a_{0}=1$ is the scale factor at the present time, $T_{0}=2.725~{}{\rm K}$ is the current temperature of CMB radiation, $\rho_{e}\equiv\rho_{\phi_{e}}$ is the energy density of the inflaton field at the end of inflation, $H_{\ast}$ is the Hubble parameter at the horizon crossing time and $N$ defines the duration of inflation from the end of inflation until the horizon passing moment . Using the conversion factor $1~{}{\rm Mpc^{-1}}=6.39\times 10^{-39}{\rm GeV}$, Eq. (41) can be simplified as follows
$N_{\rm re}=\frac{4}{1-3\omega_{\rm re}}\left[61.6-\ln\left(\frac{\rho_{e}^{1/4}}{H_{\ast}}\right)-N\right],$ (42)
@@
Here, ${\cal P}_{\rm s}(k_{\ast})=2.1\times 10^{-9}$ denotes the amplitude of the scalar power spectrum at the CMB pivot scale. It is known that, the equation of state parameter is defined as
$\omega_{\phi}\equiv\frac{p_{\phi}}{\rho_{\phi}}=-1-\frac{2}{3}\frac{\dot{H}}{H^{2}}=-1+\frac{2}{3}\epsilon.$ (44)
The value of the equation of state parameter at the end of inflation ($\epsilon=1$) is $\omega_{\phi_{e}}=-\frac{1}{3}$, and consequently
$p_{e}=-\frac{1}{3}\rho_{e},$ (45)
where $p_{e}\equiv p_{\phi_{e}}$. By substituting Eqs. (3) and (4) into Eq. (45), the energy density at the end of inflation $\rho_{e}$ in the non-canonical framework with power-law Lagrangian (2) can be obtained by
$\rho_{e}=\left(\frac{3\alpha}{\alpha+1}\right)V_{e},$ (46)
@@
For non-canonical action with the power-law Lagrangian (2), the equation of state parameter in the reheating epoch can be calculated as follows Unnikrishnan:2012
@@
Similar to the standard inflation in the canonical framework, we assume that the timescale of the scalar field oscillation is much smaller than that of the expansion of the universe. In this limit, the time variations of $\rho_{\phi}$ during each cycle is small enough to approximate $\rho_{\phi}\simeq V(\phi_{m})$, where $\phi_{m}$ is the maximum value of the field during a given cycle. According to Eq. (47), in the framework of non-canonical inflation with a power-law Lagrangian for the potential $V\left(\phi\right)=V_{0}\phi^{n}$, the equation of state parameter in the reheating epoch can be calculated as follows
$\omega_{\rm re}=\left\langle\omega_{\phi}\right\rangle=\frac{n-2\alpha}{n\left(2\alpha-1\right)+2\alpha},$ (48)
@@
$\omega_{\rm re}=\frac{2-\alpha}{5\alpha-2}.$ (49)
In the following, we calculate the duration of the reheating epoch $N_{\rm re}$ for non-canonical action with power-law Lagrangian (2) and quartic potential (35) by inserting Eqs. (43), (46) and (49) in (42). In Fig. 5, we plot the reheating duration $N_{\rm re}$ versus $\alpha$ for fixed values of the $e$-folds number $N$, as listed in Table 1. Here, a clear dependence of $N_{\rm re}$ on $\alpha$ is evident. Therefore, in the ($N_{\rm re}-\alpha$) plane, the degeneracy with respect to $\alpha$ can be broken, as different values of $\alpha$ yield distinguishable reheating durations. It is worth noting that the gray area corresponds to negative values of $N_{\rm re}$, which are non-physical and therefore excluded. In this regard, the condition $N_{\rm re}\geq 0$ imposes an upper bound on the duration of the inflationary epoch in the present model as $N=55.7$ and subsequently refine the upper bound on the non-canonical parameter to $\alpha=65$, as shown in Table 1.
Figure 5: Duration of the reheating stage $N_{\rm re}$ versus the $\alpha$ parameter for different values of the $e$-folds number $N$ according to Table 1. The gray forbidden area corresponds to negative values of $N_{\rm re}$.
@@
In addition to $N_{\rm re}$, the reheating temperature $T_{\rm re}$ can also be used to survey inflationary models and it can be calculated as follows Martin2015 ; Munoz2015 ; Cook2015 ; German2020
$T_{\rm re}=\left(\frac{30\rho_{\rm e}}{\pi^{2}g_{\rm re}}\right)^{1/4}\exp\left[-\frac{3}{4}(1+\omega_{\rm re})N_{\rm re}\right].$ (50)
@@
This parameter depends on $\omega_{\rm re}$ and $N_{\rm re}$. Since the reheating stage occurs after the end of inflation and before the beginning of the Big Bang Nucleosynthesis (BBN) stage, the upper and lower limits on the reheating temperature are defined as follows Martin2015
$10\,{\rm MeV}\leq T_{\rm re}\leq 5\times 10^{15}\,{\rm GeV}.$ (51)
Figure 6 shows the variation of $T_{\rm re}$ as a function of $\alpha$ for different fixed values of the $e$-folds number $N$, according to Table 1. The horizontal line at the top of this figure indicates the upper limit of the reheating temperature. In this case, a clear dependence on $\alpha$ is visible, and thus, in the ($T_{\rm re}-\alpha$) plane, the degeneracy can be broken.
@@
Figure 6: Reheating temperature $T_{\rm re}$ versus the parameter $\alpha$ for different values of the $e$-folds number $N$, as listed in Table 1. The maximum allowable reheating temperature, $T_{\rm re}=5\times 10^{15}~{}\rm{GeV}$, is indicated by the horizontal dashed black line at the top.
@@ IV The relic gravitational waves
Inflation theory predicts that, primordial tensor fluctuations of the inflaton field can lead to production of gravitational waves (GWs). During inflation, these fluctuations extend beyond the Hubble horizon, and upon reentering, they generate relic GWs Starobinsky:1979 ; allen88 ; sahni:1990 ; sami2002 ; dany_18 ; dany_19 ; Bernal:2019lpc . Due to their weak interaction with matter and radiation, these relic GWs serve as a valuable tool for probing the early universe. The present frequency of the originated relic GWs from the cosmic inflation can be calculated in terms of temperature using the following equation sahni:1990 ; dany_19 ; sami2002
@@
where $g_{0}^{s}=3.94$ and $g_{T}^{s}=106.75$ are the effective number of relativistic degrees of freedom in entropy at the present time and each epoch with temperature $T$, respectively. Furthermore, $g_{T}=106.75$ is the effective number of relativistic degrees of freedom in energy. The allowed ranges for the reheating temperature $T_{\rm re}$ and the corresponding frequency of GWs in the reheating epoch $f_{\rm re}$ are listed in Table 2 for several values of $\alpha$ parameter according to Table 1. The present density parameter spectrum of relic GWs in the radiation-dominated and reheating periods are calculated from the following two equations (for more information, see Ref. sahni:1990 ; dany_19 ; sami2002 )
@@
where $\Omega_{r_{0}}=2.47\times 10^{-5}h^{-2}$ is the radiation density parameter at present time, $f_{*}=2.4\times 10^{-16}~{}({\rm Hz})$ is the frequency of GWs at the CMB pivot scale $k_{\ast}=0.05~{}{\rm Mpc^{-1}}$, $f_{\rm eq}=1.7\times 10^{-17}~{}({\rm Hz})$ represents the frequency of GWs at the matter-radiation equality and $f_{e}=4.3\times 10^{8}~{}({\rm Hz})$ is the frequency of GWs at the end of inflation.
Table 2: The allowed values of reheating temperature $T_{\rm re}$ and the corresponding frequency $f_{\rm re}$ of GWs in the reheating epoch computed by Eq. (52) for several values of $\alpha$. See also Fig. 7.
$N$   $\alpha$   $T_{\rm re}/{\rm GeV}$   $f_{\rm re}/{\rm Hz}$
55   $110$   $6.29\times 10^{14}$   $2.29\times 10^{6}$
  $130$   $6.80\times 10^{14}$   $1.81\times 10^{7}$
55.3   $80$   $8.52\times 10^{14}$   $2.27\times 10^{7}$
  $130$   $1.07\times 10^{15}$   $2.84\times 10^{7}$
55.7   $60$   $1.36\times 10^{15}$   $3.63\times 10^{7}$
  $65$   $1.41\times 10^{15}$   $3.77\times 10^{7}$
Figure 7 shows the present density parameter spectra of GWs for different values of the inflationary $e$-folds number $N$, calculated using Eqs. (52)-(54). The green, red, and purple shaded regions represent the sensitivity ranges of the Big Bang Observer (BBO) Yagi:2011BBODECIGO ; Yagi:2017BBODECIGO ; Harry:2006BBO ; Crowder:2005BBO ; Corbin:2006BBO , the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) Yagi:2011BBODECIGO ; Yagi:2017BBODECIGO ; Seto:2001DECIGO ; Kawamura:2006DECIGO ; Kawamura:2011DECIGO , and the Square Kilometre Array (SKA) ska ; skaCarilli:2004 ; skaWeltman:2020 , respectively.
The reheating temperatures and gravitational wave frequencies at the breakpoints of the spectra are listed in Table 2. In each graph, the blue and red curves represent the present energy density of GWs, corresponding to the maximum and minimum allowed values of $\alpha$ for each $e$-folds number $N$, respectively.
It can be observed that, the curves fall within the sensitivity range of GWs detectors. Hence, with the advancement of these detectors, there is a possibility of detecting the predicted relic GWs by this model. Moreover, varying in $\alpha$ parameter leads to distinct values of relic GWs density.
Figure 7: The energy density spectrum of relic GWs as a function of frequency for various $e$-folds numbers (a) $N=55$, (b) $N=55.3$, and (c) $N=55.7$, corresponding to different values of $\alpha$. The shaded regions represent the sensitivity ranges of GW detectors, including BBO Yagi:2011BBODECIGO ; Yagi:2017BBODECIGO ; Harry:2006BBO ; Crowder:2005BBO ; Corbin:2006BBO , DECIGO Yagi:2011BBODECIGO ; Yagi:2017BBODECIGO ; Seto:2001DECIGO ; Kawamura:2006DECIGO ; Kawamura:2011DECIGO , SKA ska ; skaCarilli:2004 ; skaWeltman:2020 . In each graph, the blue (red) curve corresponds to the minimum (maximum) allowed value of $\alpha$, as given in Table 1. The break in each spectrum marks the reheating frequency $f_{\rm re}$, which is related to the reheating temperature $T_{\rm re}$ listed in Table 2.